OpenStudy (anonymous):

prove that arctan(x)=1-x has at least one root

OpenStudy (anonymous):

it is continuous for sure, so show that it is negative somewhere and positive somewhere else. try say \[-\frac{\pi}{4}\] and \[\frac{\pi}{3}\] so make your life easy

OpenStudy (rogue):

That's utilizing the intermediate value theorem right?

OpenStudy (anonymous):

Yes